A gas exerts a pressure of 2.0 atm, at 30°C, in a 10 L container. In what size container would the same amount of gas exert a pressure of 4.0 atm at 20° C?

  


1
Expert's answer
2021-12-20T09:17:02-0500

Using the gas law:

pV = nRT, and keeping the same number of moles of gas (n),

we find the final volume V2:

V2 = (T2/p2)*(p1*V1/T1) = (293 K/ 4.0 atm)*(2.0 atm * 10 L / 303 K) = 4.83 L